# Majorization via positivity of Jack and Macdonald polynomial differences

### v1

@@ 4.2. Jack polynomial differences with varying parameter
Since $M_{\lambda}(x)=P_{\lambda}(x;0)/P_{\lambda}(\bm{1};0)$, Eq. 4.1 can be viewed as a positivity property involving Jack polynomial differences with different parameters. Let us generalize this idea further.
@@
Fix $0\leqslant\sigma\leqslant\tau\leqslant\infty$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference with parameter $\tau$ is positive in the Jack semiring $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$:
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0}).$ (4.3)
Note that Conjecture 4 is false when $\tau<\sigma$. For example, let $n=3$, $\tau=0$, $\sigma=1$, then $M_{(3)}-M_{(21)}=\frac{10}{3}S_{(3)}-4S_{(21)}+\frac{2}{3}S_{(111)}$, where $S_{\lambda}=s_{\lambda}/s_{\lambda}(\bm{1})$.
Also note that the expansion of $P_{\lambda}(x;\tau)$ into a sum of polynomials $P_{\mu}(x;\sigma)$ with $\sigma\leqslant\tau$ may not be positive. For example, when $n=3$, we have (suppressing the $x$)
$P_{(33)}(\tau)=P_{(33)}(\sigma)+\frac{6(\tau-\sigma)}{(2+\sigma)(2+\tau)}P_{(321)}(\sigma)+\frac{6(\tau-2\sigma)(\tau-\sigma)}{(1+\sigma)(1+2\sigma)(1+\tau)(2+\tau)}P_{(222)}(\sigma),$
and so the last coefficient is negative if $\sigma<\tau<2\sigma$.
When $\sigma=1$, these coefficients are falling factorials in $\tau$. There is a combinatorial conjecture in this case in [2]. (It should be noted that the parameter $\alpha$ in [2] is not the same as Macdonald’s $\alpha$, instead, it is the same as our $\tau$.)
Now, we again look at the case $n=2$. This time, the computation is more complicated.
Example 4.7 (Toy Example for Conjecture 4).
Let $n=2$ and $d\geqslant 2$. In this example, we mostly suppress the argument $x$, and also assume the indices $i,j=0,\dots,\lfloor d/2\rfloor$ unless otherwise stated.
We prove the stronger statement the difference in Eq. 4.3 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. By Sections 3.2 and 3.2, we have
@@
$M_{d}(P(\tau),m)\coloneqq\left(\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\right).$
@@
$M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)\coloneqq\left(\frac{1}{P_{(d-i,i)}(\bm{1};\tau)}\delta_{ij}\right)=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-i}}}\delta_{ij}\right).$
We claim that its inverse is given by
$M_{d}\left(P(\sigma),m\right)^{-1}=\left((-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}\frac{(\sigma+i-j+1)_{{d-i-j}}(\sigma+d-i-j+1)_{{j-i}}}{(\sigma+1)_{{d-2i}}}\right).$
(When $i=j=d/2$, the factor $\frac{d-2i}{d-i-j}$ is understood to be 1.) For example, if $d=2,3$, then
$M_{2}(P(\tau),m)=\begin{pmatrix}1&-\frac{2\tau}{1+\tau}\\ &1\end{pmatrix},\quad M_{3}(P(\tau),m)=\begin{pmatrix}1&-\frac{3\tau}{1+\tau}\\ &1\end{pmatrix}$,
Note that $(-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}$ is related to the Cardan polynomial $2T_{n}(x/2)$ (where $T_{n}(x)$ is the Chebyshev polynomial), see [OEIS A034807].
We claim that the transition matrix $M_{d}\left(P(\tau),P(\sigma)\right)$ between Jack polynomials with different parameters, which is given by the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$, can be computed by
$M_{d}\left(P(\tau),P(\sigma)\right)$ $=M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$
$=\left(\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\right).$ (4.4)
(When $i>j$, the entries are understood to be 0.) Note that the first claim follows from the second, since upon setting $\tau=\sigma$, Section 4.2 yields the identity matrix.
To prove Section 4.2, it suffices to check the strictly upper diagonal entries. Let $i<j$, and compute the $(i,j)$-entry of the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$:
@@
Denote by $f(i,j,d;l)$ the summand on the left, and $r(i,j,d)$ the RHS. We shall prove Section 4.2 using the so-called Wilf–Zeilberger algorithm [23, Chapter 7]. Let
@@
and so for any $j>i$, $\sum_{l=0}^{j-i}F(i,j,d;l)=\sum_{l\geqslant 0}F(i,j,d;l)=1$. We have thus proved Section 4.2. The transition matrix from $\left(\frac{P_{\lambda}(\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $\left(\frac{P_{\lambda}(\sigma)}{P_{\lambda}(\bm{1};\sigma)}\right)$ is
@@
where $\propto$ means that a positive factor is omitted. It suffices to show that $a_{j}<0$ for $j=j_{0},\dots,\lfloor d/2\rfloor$ for some $j_{0}\geqslant 1$.
@@
Also, we have $a_{j+1}-a_{j}=-2(2\tau-1)(\sigma+d-2j-1)$. If $\tau\geqslant\frac{1}{2}$, $(a_{j})$ is decreasing, in particular, $a_{j}$ satisfies the desired condition. If $\tau<\frac{1}{2}$, $(a_{j})$ is increasing, hence $a_{j}\leqslant a_{\lfloor d/2\rfloor}<0$.
@@
Again, we prove the stronger statement the difference in Eq. 4.3 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. First, let $n=2$. The example above shows that Conjecture 4 holds for the pair $\lambda=(d,0)\mathrel{\succcurlyeq}\mu=(d-1,1)$. It suffices to show that it holds for any adjacent pair $\lambda=(d-i,i)\mathrel{\succcurlyeq}\mu=(d-i-1,i+1)$, as the general case would follow by telescoping. But this holds because
@@
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu}(x;\tau)}{P_{\nu}(\bm{1};\tau)}-\frac{P_{\xi}(x;\tau)}{P_{\xi}(\bm{1};\tau)}\right)$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu+(i,i)}(x;\tau)}{P_{\nu+(i,i)}(\bm{1};\tau)}-\frac{P_{\xi+(i,i)}(x;\tau)}{P_{\xi+(i,i)}(\bm{1};\tau)}\right).$
@@
4.3. A conjecture of Alexandersson–Haglund–Wang
For completeness, we end this section by recalling a related conjecture in [2]. Following the convention in [4, Section 6.2], we define the integral Jack polynomials as
$J_{\lambda}(x;\tau)=c_{\lambda}(\tau)P_{\lambda}(x;\tau),\quad\text{where}\quad c_{\lambda}(\tau)\coloneqq\prod_{s\in\lambda}(a_{\lambda}(s)+\tau(l_{\lambda}(s)+1)).$ (4.7)
For $s=(i,j)\in\lambda$, we have $a_{\lambda}(s)\coloneqq\lambda_{i}-j$ and $l_{\lambda}(s)=\lambda_{j}^{\prime}-i$ are the number of boxes direct to the East and South of $s$ (drawn in the English manner).
Note that our $J_{\lambda}(x;\tau)$ is the same as the $\tilde{J}_{\lambda}^{(\alpha)}(x)$ in [2], with $\alpha$ replaced by $\tau$ (note: not $\frac{1}{\tau}$); and is related to the $J_{\lambda}^{(\alpha)}(x)$ in [18, VI. (10.22)] by $J_{\lambda}(x;\tau)=\tau^{-|\lambda|}J^{(1/\tau)}_{\lambda}(x)$.
As shown in [12], the expansion coefficients of integral Jack into the (augmented) monomial basis are positive and integral and can be combinatorially interpreted as admissible (also known as non-attacking) tableaux.
It was recently conjectured in [2] that the expansion coefficients of the integral Jack $J_{\lambda}(x;\tau)$ into the Schur basis $s_{\mu}$ have certain positivity and integrality. Define
$v_{\lambda\mu}(\tau)\coloneqq\langle J_{\lambda}(\tau),s_{\mu}\rangle,\qquad\text{that is,}\qquad J_{\lambda}(\tau)=\sum_{\mu}v_{\lambda\mu}(\tau)s_{\mu}.$ (4.8)
Conjecture 5 ([2]).
Let $\lambda,\mu$ be partitions of $d$, and define the expansion coefficients $a_{k}(\lambda,\mu)$ and $b_{k}(\lambda,\mu)$ as follows:
$v_{\lambda\mu}(\tau)=\sum_{k=0}^{d-1}a_{k}(\lambda,\mu)\binom{\tau+k}{d}=\sum_{k=1}^{d}b_{d-k}(\lambda,\mu)\binom{\tau}{k}k!.$ (4.9)
Then $a_{k}(\lambda,\mu)$ and $b_{d-k}(\lambda,\mu)$ are in $\mathbb{Z}_{\geqslant 0}$. Furthermore, the polynomials $\sum_{k=0}^{d}a_{k}(\lambda,\mu)z^{k}$ and $\sum_{k=0}^{d}b_{d-k}(\lambda,\mu)z^{k}$ have only real zeros.
@@ 5. Macdonald polynomial differences: conjectures and results
Let $P_{\lambda}(x)=P_{\lambda}(x;q,t)$ denote the monic Macdonald polynomials. Our goal in this section is to explore if Conjecture 3, characterizing majorization in terms of (normalized) Jack differences belonging to the Muirhead semiring, can be strengthened to Macdonald polynomials – leading to a strengthening of Conjecture 1 as well. And indeed, we achieve both of these goals, in the cases where we have shown the conjectures above.
@@
We now study Macdonald differences for $\lambda\succcurlyeq\mu$, starting with the case of two variables. The first question in extending the aforementioned conjectures is: how should one normalize the Macdonald polynomials. We begin by exploring the natural choice [18] of $x=t^{\delta}=(t^{n-1},\dots,t,1)$.
@@
Let $n=2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. We have, using the combinatorial formula [18, VI. (7.13)],
@@
By [18, VI. (6.11)], where $n=2$ and $t^{\delta}=(t,1)$, $P_{\lambda}(t^{\delta})=\frac{(t^{2};q)_{{d}}}{(t;q)_{{d}}}$. Hence,
@@
$P_{\lambda}(\bm{1})=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}.$
@@
$\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}\text{\quad and \quad}\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}=\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}.$
@@
$\mathrel{\phantom{=}}\frac{P_{\lambda}(x)}{P_{\lambda}(\bm{1})}-\frac{P_{\mu}(x)}{P_{\mu}(\bm{1})}$
@@
$b_{i+1}-b_{i}=\dfrac{1-tq^{i}}{1-q^{i+1}}-\dfrac{1-tq^{i-1}}{1-q^{i}}=\frac{q^{i-1}(q-t)(q-1)}{(1-q^{i})(1-q^{i+1})}.$
Hence $(b_{i})_{i}$ is increasing if $q>t$ and decreasing if $q<t$.
(2) We first show that the two expressions are equal. Separate the positive term for $i=l$, and collect the negative term for $i$ and the positive term for i1i-1; now the last term above becomes the positive term for $i=1$. If $q=t$, the $(2l+1)$-term polynomial becomes $1$. If $q>t$, then by part (1), $b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}}>0$. If $q<t$, then
@@
As the expression within each parentheses is of the form in Eq. 5.4, it is positive. ∎
Thus, we arrive at the following conjecture for normalized Macdonald differences in the Muirhead semiring – and we now merge this with the Macdonald analogue of the Cuttler–Greene–Skandera conjecture too. Recall here that $\mathbb{F}_{\geqslant 0}$ and $\mathbb{F}_{\geqslant 0}^{\mathbb{R}}$ for the Macdonald case are defined in Eq. 2.12.
@@
Suppose $\lambda$ and $\mu$ are partitions with $|\lambda|=|\mu|$. Then the following are equivalent:
@@
    The Macdonald difference lies in the Muirhead semiring over $\mathbb{F}_{\geqslant 0}$:
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{F}_{\geqslant 0}).$ (5.5)
@@
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.6)

### v2

@@ 4.2. Jack polynomial differences with varying parameter
Since $m_{\lambda}(x)=P_{\lambda}(x;\tau=0)$, Eq. 4.3 can be viewed as a positivity property involving Jack polynomial differences in two different parameters $\tau$ and 0. Let us generalize this idea further.
@@
Fix $0\leqslant\sigma<\tau\leqslant\infty$. If $\lambda\mathrel{\succcurlyeq}\mu$, then the Jack difference with parameter $\tau$ is positive in the Jack semiring $\mathcal{J}_{S}^{\sigma}(\mathbb{R}_{\geqslant 0})$:
$\frac{P_{\lambda}(x;\tau)}{P_{\lambda}(\bm{1};\tau)}-\frac{P_{\mu}(x;\tau)}{P_{\mu}(\bm{1};\tau)}\in\mathcal{J}_{S}^{\sigma}(\mathbb{R}_{\geqslant 0}).$ (4.5)
The following example shows that Conjecture 4 is false when $\tau<\sigma$.
Example 4.7.
Consider $n=3$, $\tau=0$, $\sigma=1$, and let
$F\coloneqq 3(M_{(3)}-M_{(21)})=10S_{(3)}-12S_{(21)}+2S_{(111)},$
where $S_{\lambda}=s_{\lambda}/s_{\lambda}(\bm{1})$. Clearly, it does not belong to the Schur cone. We claim that $F$ does not belong to the Schur semiring.
Note that in degree 3, the Schur semiring is generated by
$f_{1}$ $\coloneqq S_{(1)}^{3},\quad f_{2}\coloneqq S_{(1)}S_{(11)},\quad f_{3}\coloneqq S_{(111)},$
$f_{4}$ $\coloneqq S_{(1)}\bigl(S_{(2)}-S_{(11)}\bigr),\quad f_{5}\coloneqq S_{(21)}-S_{(111)},\quad f_{6}\coloneqq S_{(3)}-S_{(21)}.$
It is not hard to see that $f_{3},f_{5},f_{6}$ are enough to generate the semiring since
$f_{1}=f_{3}+\frac{26}{27}\,f_{5}+\frac{10}{27}\,f_{6},\quad f_{2}=f_{3}+\frac{8}{9}\,f_{5},\quad f_{4}=\frac{1}{9}\,f_{5}+\frac{5}{9}\,f_{6}.$
However, $F=10f_{6}-2f_{5}$, hence $F$ does not lie in the Schur semiring.
Now, we again look at the two-variable case in Section 3.3. This time, the computation is more complicated.
Example 4.8.
Consider $n=2$, $d\geqslant 2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. In this example, we mostly suppress the argument $x$, and also assume the indices $i,j=0,\dots,\lfloor d/2\rfloor$ unless otherwise stated.
We prove the stronger statement the difference in Eq. 4.5 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$.
$M_{d}\left(\frac{P(\tau)}{P(\bm{1};\tau)},P(\tau)\right)\coloneqq\left(\frac{1}{P_{(d-i,i)}(\bm{1};\tau)}\delta_{ij}\right)_{i,j}=\left(\frac{(\tau)_{{d-2i}}}{(2\tau)_{{d-i}}}\delta_{ij}\right)_{i,j}.$ (4.6)
By Sections 3.2.1 and 3.2.1, we have
@@
$M_{d}(P(\tau),m)\coloneqq\left(\binom{d-2i}{j-i}\frac{(\tau)_{{j-i}}(\tau)_{{d-j-i}}}{(\tau)_{{d-2i}}}\right)_{i,j}.$
@@
Define a matrix
$N_{d}(\sigma)=\left((-1)^{i-j}\binom{d-i-j}{j-i}\frac{d-2i}{d-i-j}\frac{(\sigma+i-j+1)_{{d-i-j}}(\sigma+d-i-j+1)_{{j-i}}}{(\sigma+1)_{{d-2i}}}\right)_{i,j}.$ (4.7)
(When $i=j=d/2$, the factor $\frac{d-2i}{d-i-j}$ is understood to be 1.) For example, when $d=2,3$, we have
$N_{d}(\sigma)=\begin{pmatrix}1&-\frac{2\sigma}{1+\sigma}\\ &1\end{pmatrix},\quad N_{d}(\sigma)=\begin{pmatrix}1&-\frac{3\sigma}{1+\sigma}\\ &1\end{pmatrix}$.
We claim that the product matrix $M_{d}\left(P(\tau),m\right)N_{d}(\sigma)$ is given by
$M_{d}\left(P(\tau),m\right)N_{d}(\sigma)$ $=\left(\frac{(d-2i)!}{(d-2j)!(j-i)!}\frac{(\tau-\sigma)_{{j-i}}}{(\tau+d-i-j)_{{j-i}}(\sigma+d-2j+1)_{{j-i}}}\right)_{i,j}.$ (4.8)
(When $i>j$, the entries are understood to be 0.) In particular, when setting $\sigma=\tau$, we see that the product $M_{d}\left(P(\tau),m\right)N_{d}(\tau)$ is the identity matrix, and hence $N_{d}(\tau)=M_{d}\left(P(\tau),m\right)^{-1}$ is the inverse transition matrix, and $M_{d}\left(P(\tau),m\right)N_{d}(\sigma)=M_{d}\left(P(\tau),P(\sigma)\right)$ is the transition matrix between Jack polynomials with different parameters.
Given Eq. 4.7, to prove Eq. 4.8, it suffices to check the strictly upper diagonal entries. Let $i<j$, and compute the $(i,j)$-entry of the product $M_{d}\left(P(\tau),m\right)M_{d}\left(P(\sigma),m\right)^{-1}$:
@@
Denote by $f(i,j,d;l)$ the summand on the left, and $r(i,j,d)$ the RHS. We prove Example 4.8 using the Wilf–Zeilberger algorithm [22, Chapter 7]. Let
@@
and so for any $j>i$, $\sum_{l=0}^{j-i}F(i,j,d;l)=\sum_{l\geqslant 0}F(i,j,d;l)=1$. We have thus proved Eq. 4.8.
Next, we show (as asserted just before Eq. 4.6 that the difference in Eq. 4.5 belongs to the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. The transition matrix from $\left(\frac{P_{\lambda}(\tau)}{P_{\lambda}(\bm{1};\tau)}\right)$ to $\left(\frac{P_{\lambda}(\sigma)}{P_{\lambda}(\bm{1};\sigma)}\right)$ is
@@
where $\propto$ means that a factor in $\mathbb{F}_{\geqslant 0}$ has been omitted.
@@
For $j=1,\dots,\lfloor d/2\rfloor-1$, we have $a_{j+1}-a_{j}=-2(2\tau-1)(\sigma+d-2j-1)$, hence $(a_{j})$ is monotone.
Now we have the expansion
$\frac{P_{(d,0)}(\tau)}{P_{(d,0)}(\bm{1};\tau)}-\frac{P_{(d-1,1)}(\tau)}{P_{(d-1,1)}(\bm{1};\tau)}=\frac{P_{(d,0)}(\bm{1};\sigma)}{P_{(d,0)}(\bm{1};\tau)}\frac{P_{(d-j,j)}(\sigma)}{P_{(d-j,j)}(\bm{1};\sigma)}+\sum_{j=1}^{\lfloor d/2\rfloor}c_{j}a_{j}\frac{P_{(d-j,j)}(\sigma)}{P_{(d-j,j)}(\bm{1};\sigma)},$
where $c_{j}>0$ denotes the factor omitted above, and $(a_{j})$ is monotone and ends at a negative number. In particular, there exists some $j_{0}$ such that $a_{j}\geqslant 0$ for $j<j_{0}$ and $a_{j}\leqslant 0$ for $j\geqslant j_{0}$. Such an expansion lies in the Jack cone by Abel’s lemma.
@@
Again, we prove the stronger statement the difference in Eq. 4.5 lies in the Jack cone $\mathcal{J}_{C}^{\sigma}(\mathbb{R}_{\geqslant 0})$. First, let $n=2$. The example above shows that Conjecture 4 holds for the pair $\lambda=(d,0)\mathrel{\succcurlyeq}\mu=(d-1,1)$. It suffices to show that it holds for any adjacent pair $\lambda=(d-i,i)\mathrel{\succcurlyeq}\mu=(d-i-1,i+1)$, as the general case would follow by telescoping. This holds because
@@
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu}(x;\sigma)}{P_{\nu}(\bm{1};\sigma)}-\frac{P_{\xi}(x;\sigma)}{P_{\xi}(\bm{1};\sigma)}\right)$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{R}_{\geqslant 0}\cdot\left(\frac{P_{\nu+(i,i)}(x;\sigma)}{P_{\nu+(i,i)}(\bm{1};\sigma)}-\frac{P_{\xi+(i,i)}(x;\sigma)}{P_{\xi+(i,i)}(\bm{1};\sigma)}\right)\subseteq\mathcal{J}_{C}(\mathbb{R}_{\geqslant 0}).$
@@ 5. Macdonald polynomial differences: conjectures and results
Let $P_{\lambda}(x)=P_{\lambda}(x;q,t)$ denote the monic Macdonald polynomials. Our goal in this section is to explore if Conjecture 3, characterizing majorization in terms of (normalized) Jack differences belonging to the Muirhead semiring, can be strengthened to Macdonald polynomials—leading to a strengthening of Conjecture 1 as well. And indeed, we achieve both of these goals, in the cases where we have shown the conjectures above.
@@
We now study Macdonald differences for $\lambda\mathrel{\succcurlyeq}\mu$, starting with the case of two variables. The first question in extending the aforementioned conjectures is: how should one normalize the Macdonald polynomials. We begin by exploring the natural choice [17] of $x=t^{\delta}=(t^{n-1},\dots,t,1)$.
@@
Let $n=2$, $\lambda=(d,0)$ and $\mu=(d-1,1)$. We have, using the combinatorial formula [17, VI. (7.13)],
@@
By [17, VI. (6.11)], where $n=2$ and $t^{\delta}=(t,1)$, $P_{\lambda}(t^{\delta})=\frac{(t^{2};q)_{{d}}}{(t;q)_{{d}}}$. Hence,
@@
$P_{\lambda}(1,1)=\sum_{i=0}^{d}\frac{(q;q)_{{d}}}{(q;q)_{{i}}(q;q)_{{d-i}}}\frac{(t;q)_{{i}}(t;q)_{{d-i}}}{(t;q)_{{d}}}=\frac{(q;q)_{{d}}}{(t;q)_{{d}}}\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}.$
@@
$\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}=\frac{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=0}^{d}\frac{(t;q)_{{i}}}{(q;q)_{{i}}}\frac{(t;q)_{{d-i}}}{(q;q)_{{d-i}}}}\text{\quad and \quad}\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}=\frac{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}x_{1}^{i}x_{2}^{d-i}}{\displaystyle\sum_{i=1}^{d-1}\frac{(t;q)_{{i-1}}}{(q;q)_{{i-1}}}\frac{(t;q)_{{d-1-i}}}{(q;q)_{{d-1-i}}}}.$
@@
$\mathrel{\phantom{=}}\frac{P_{\lambda}(x_{1},x_{2})}{P_{\lambda}(1,1)}-\frac{P_{\mu}(x_{1},x_{2})}{P_{\mu}(1,1)}$
@@
$b_{i+1}-b_{i}=\dfrac{1-tq^{i}}{1-q^{i+1}}-\dfrac{1-tq^{i-1}}{1-q^{i}}=\frac{q^{i-1}(t-q)(1-q)}{(1-q^{i})(1-q^{i+1})}.$
Hence $(b_{i})$ is increasing if $q<t$ and decreasing if $q>t$.
(2) We first show that the two expressions are equal. Separate the positive term for $i=l$, and collect the negative term for $i$ and the positive term for i1i-1; now the last term above becomes the positive term for $i=1$. If $q=t$, the $(2l+1)$-term polynomial becomes $1$. If $q<t$, then by part (1), $b_{d-i}b_{d-i+1}-b_{s_{i}}b_{d-s_{i}}>0$. If $q>t$, then
@@
As the expression within each parentheses is of the form in Eq. 5.4, it is positive.
Thus, we arrive at the following conjecture for normalized Macdonald differences in the Muirhead semiring—and we now merge this with the Macdonald analogue of the Cuttler–Greene–Skandera conjecture too. Let $\mathbb{K}_{\geqslant 0}$ and $\mathbb{K}_{\geqslant 0}^{\mathbb{R}}$ be the cones for the Macdonald case, as defined in Eq. 2.12.
@@
The following are equivalent for partitions $\lambda$ and $\mu$ :
@@
    The Macdonald difference lies in the Muirhead semiring over $\mathbb{K}_{\geqslant 0}$:
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathcal{M}_{S}(\mathbb{K}_{\geqslant 0}).$ (5.5)
@@
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{K}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.6)
  3. $(2^{\prime})$
    For every $q_{0},t_{0}\in(0,1)$, the Macdonald difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$: